OpenStudy (kaylardz0405):

Simplify the expression. –16 • 1/8 • 4 + 32 ÷ 2 A. –36 B. 24 C. 8 D. 12

OpenStudy (kaylardz0405):

will fan and medal

OpenStudy (kaylardz0405):

is it b

jagr2713 (jagr2713):

Hey B is wrong D: what do you think we do first?

OpenStudy (kaylardz0405):

multiply -16 and 1/8

jagr2713 (jagr2713):

what do we get?

OpenStudy (kaylardz0405):

-2

jagr2713 (jagr2713):

Good but lets keep it as a fraction so we get \[-\frac{ 16 }{ 8 }\]

OpenStudy (kaylardz0405):

ok

jagr2713 (jagr2713):

so we have \[-\frac{ 16 }{ 8 }\times 4\] what do we get?

OpenStudy (kaylardz0405):

64/8

jagr2713 (jagr2713):

you forgot some D:

jagr2713 (jagr2713):

-16/8 X 4

OpenStudy (kaylardz0405):

- 64/8

jagr2713 (jagr2713):

Good we we have \[-\frac{ 64 }{ 8 }+\frac{ 32 }{ 2 }\] what do we simplify next?

OpenStudy (kaylardz0405):

96/10

jagr2713 (jagr2713):

we simplify -64/8 what do we get? and what are we left with?

OpenStudy (kaylardz0405):

ok i got it thanks

jagr2713 (jagr2713):

What did u get?

jagr2713 (jagr2713):

Ok we get -8 and we have\[-8+\frac{ 32 }{ 2 }\] Then we simplify 32/2 to 16 and left with -8+16 so now we add and get 8 as the answer :D Hope you got the same thing and\[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color {green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid} {\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed} {\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color {magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color {ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] if you want to know more about this site, Here is a tutorial about this site : http://openstudy.com/study#/updates/5525c0fce4b030ceb3f357e8